# Learning-Based Two-Way Communications: Algorithmic Framework and Comparative Analysis

### v1

@@ IV Results and Discussion
In this section, we first analytically determine the computational requirements for the encoders and decoders in terms of FLOPS. We then analyze the performance of two-way versus one-way coding under symmetric SNR conditions, and we conclude by comparing the two-way models’ efficacies across a variety of SNR regimes. We standardize the users’ message lengths by setting $K_{1}=K_{2}=K$ [^4] Simulations were conducted on an NVIDIA A100 40GB GPU. Code can be found at: https://github.com/DavidRNickel/twoway_models.git.
@@ IV-A Comparison of Floating-Point Operations (FLOPS)
We quantify the computational overhead by calculating the total number of FLOPS to process a single bitstream. Let $h_{a}$ be the size of a model’s hidden dimensions, with $a\in\{c,b,r\}$, where $c,b,r$ denote TWLC, TWBAF, and TWRNN, resp.; $\ell$ be the number of rows in a TWBAF input matrix; $L_{B}^{E}$ and $L_{B}^{D}$ be the number of attention layers in TWBAF’s encoder and decoder, resp.; and $K,~{}M,~{}T$ be the block-length, sub-block-length, and number of channel uses, resp., with $T_{M}=TM/K$.
We report the total FLOPS for each method in Tab. I. For comparison, consider setting $K=6$, and $T=18$. For TWLC set $h_{c}=32$, $M=3$; for TWRNN, set $h_{r}=50$, $M=3$; and for TWBAF, set $h_{b}=32$, $M=2$, $\ell=3$, $L_{B}^{E}=2$, and $L_{B}^{D}=3$. Then the total FLOPS counts are $0.598M$, $2.53M$, and $2.61M$ per Tab. I, for TWLC, TWBAF, and TWRNN, respectively. The training time of TWLC, TWRNN, and TWBAF totaled approximately 3, 9, and 24 hours, respectively. The relatively small size of TWLC enabled it to train more efficiently than the other two on a large batch size of $10^{5}$ (per user), whereas TWRNN and TWBAF were forced to train on $2.5\cdot 10^{4}$ and $10^{4}$, respectively, so as to fit onto the GPU. Lastly, we note that the potential parallelizability of TWBAF’s attention operations may afford it an advantage over TWRNN in regimes such as $SNR_{1}=1~{}\text{dB},~{}SNR_{2}=15$ dB, where they obtain similar BLERs (Fig. 4).
TABLE I: FLOPS comparison between different models, where $K,~{}M,~{}T$ be the overall block-length, sub-block-length, and total number of channel uses, respectively. Equations are for a single user’s encoder and decoder, i.e., half of the total FLOPS count.
@@
TWLC $T(4h_{c}(M+2(T_{M}-1))+26h_{c}^{2}+22h_{c}+3)$ $\frac{K}{M}[4h_{c}(M\!+\!T_{M})\!+\!28h_{c}^{2}+20h_{c}+2^{M}(2h_{c}+3)-1]$
TWBAF $T_{M}[L_{B}^{E}(36\ell h_{b}^{2}\!+\!38\ell h_{b})+\ell h_{b}+36\ell(2(T_{M}-1)+M)h_{b}+36\ell h_{b}^{2}+11\ell+h_{b}\ell/2]$ $L_{B}^{D}(36\ell h_{b}^{2}+38\ell h_{b})+6\ell(M+T_{M})h_{b}+36\ell h_{b}^{2}+9\ell+2^{M+1}\ell h_{b}/4+\ell(3\cdot 2^{M}-1)$
TWRNN $T[18h_{r}^{2}+h_{r}(6M+48)+11]$ $\nicefrac{{K}}{{M}}[T_{M}(36h_{r}^{2}\!+\!18Mh_{r}+73h_{r})+2^{M}(4h_{r}+3)-1]$
@@ IV-B Comparison of One-way Versus Two-way Communications
In this section, we explore the performance of one-way versus two-way communications. Tab. II demonstrates results for $SNR_{2}=20$ dB. ALC shows considerable BLER improvement over original LC, thus making it a more competitive model through which to compare one- and two-way coding.
We now compare the performance of TWLC against independent instances of one-way ALC sharing the channel in a time-division manner. The users transmit $K$ bits apiece over a total of $T$ channel uses. For the one-way case, each user is allotted $T/2$ channel uses. For the two-way model, the bitstreams of length $K$ are split into two sub-blocks of length $M=K/2$, wherein those $M$ bits are transmitted over $T_{M}=T/2$ channel uses. The decoders’ outputs are then concatenated for calculation of BLERs. Fig. 3 sets $SNR_{1}=SNR_{2}$. For $K=4,~{}T=16$, in lower SNR regimes the two-way case bests its one-way counterpart, and vice versa for higher SNR. However, for $K=6,~{}T=18$, the lower coding rate enabled by the two-way paradigm allows the users to consistently achieve a much lower sum-BLER. While this may be in part attributable to the specfic schemes compared (i.e., TWLC and ALC), we believe that the underlying impact on the coding rate afforded by jointly encoding bits and feedback information plays a great role in determining the error rate.
Figure 3: Comparison of one-way vs two-way feedback coding with symmetric SNRs using Active LC and TWLC.
TABLE II: BLER Performance of one-way Active LightCode ($BLER(ALC)$) vs standard LightCode ($BLER(LC)$) for $SNR_{2}=20$ with $K=M=3$ and $T=9$.
@@
BLER(ALC) $45.6\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 4 $61.3\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 6 $89.8\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 8
@@ IV-C Two-Way BLER Comparison
Figure 4: Comparisons between rate $\nicefrac{{1}}{{3}}$ two-way models under different SNR settings. In (a) and (b), $K$ denotes overall block-length, and $M\leq K$ denotes the sub-block-length. In (a) with $SNR_{1}$ set to $1$ dB, all models perform reasonably well at lower $SNR_{2}$, while TWBAF and TWRNN showing obvious advantage at higher at higher $SNR_{2}$. (b) shows TWRNN with a clear advantage with $K=6,~{}M=6$. In (c), we compare the sensitivity of models trained on $SNR_{1}=1,~{}SNR_{2}=15$ (labeled “Fixed”) running on varying $SNR_{2}$. Lines marked “Orig” are the datapoints from (a), for which the training and testing SNRs are matched for the model corresponding to each datapoint.
In Fig. 4, we explore rate $1/3$ codes on messages of length $K=6$. For length-$M$ sub-blocks ($M\leq K$), $TM/K$ channel uses apiece are utilized in transmitting $K/M$ sub-blocks. The decoded results of these $K/M$ sub-block transmissions are concatenated for BLER calculation of the length-$K$ block. In Fig. 4(a), for $SNR_{1}=1$ dB, all models perform comparably in lower $SNR_{2}$ regimes, making TWLC, with its smaller architecture, a viable option. TWRNN and TWBAF perform substantially better as $SNR_{2}$ increases, indicating their ability to capture more complex relationships between the users. For $SNR_{1}=-1$ dB in Fig. 4(b), we find that TWRNN with $M=6$ clearly outperforms all other models. Lastly in Fig. 4(c), the sensitivity of each model is assessed against out-of-distribution noise by first training a “fixed” model on $SNR_{1}=1~{}\text{dB},~{}SNR_{2}=15$ dB then testing the model as a function of $SNR_{2}$. While all models are reasonably robust to noises other than those which they were trained on, there is distinct benefit in tailoring the training to closely match the intended use-case, as evidenced by the superior performance of the lines labeled “Orig”, run with matched train/test SNR.
@@ IV-D Key Takeaways
TWLC, TWBAF, and TWRNN all demonstrate viability under different engineering goals and system settings. From Sec. IV-B, we find that one-way and two-way communications demonstrate distinct SNR regions and coding rates over which they respectively perform better. Examining the distinct models, we find that TWLC obtains decent performance with a significantly smaller network, requiring under $25\%$ of the FLOPS to run as the other two (Sec. IV-A). However, TWLC is relatively less capable of capturing the intricate patterns available in certain SNR regimes. TWRNN demonstrates superior BLER performance in low SNR regimes but suffers from large size and computational cost due to its recurrent structure. Lastly, TWBAF performs comparably to TWRNN in higher $SNR_{1}$ settings, with its better parallelizability making it perhaps the more desirable choice. The high performance of TWRNN and TWBAF appears in large part to be due to their utilization of attention mechanisms, their major commonality.
@@ V Conclusions
In this paper, we presented a system model for two-way feedback coding. We then explored the performance tradeoffs of one-way and two-way communications. We find that in some noise regimes, two-way coding demonstrates superior performance to comparable one-way systems. Lastly, we explored the efficacy of three different two-way models, TWLC, TWRNN, and TWBAF, finding that each is meritorious depending on the desired engineering outcomes. These efforts contribute starting foundational benchmarks for the emerging field, offering valuable baselines to guide and compare future research endeavors, including the development of novel coding schemes, adaptive feedback mechanisms, and performance optimization under diverse channel conditions.

### v2

@@ IV Results and Discussion
In this section, we first assess the computational requirements for the encoders and decoders in terms of FLOPS. We then analyze the performance of TW versus OW coding under symmetric SNRs, and we conclude by comparing the TW models’ efficacies across a variety of SNR regimes. We set message lengths to $K_{i}=K$ and powers to $P_{i}=0$ dB, $\forall i$. For reference, we provide sum BLER results for $(T,K)$ polar codes (as in e.g., [7]) over an open-loop TW channel. [^5] Code available: https://github.com/DavidRNickel/twoway_models.git.
@@ IV-A Comparison of Floating-Point Operations (FLOPS)
We quantify the computational overhead by calculating the total number of FLOPS to process a single bitstream. Let $h_{\{c,b,r\}}$ be the size of a model’s hidden dimensions, where $c,b,r$ denote TWLC, TWBAF, and TWRNN, resp.; $\ell$ be the number of rows in a TWBAF input matrix; $L_{B}^{E}$ and $L_{B}^{D}$ be the number of attention layers in TWBAF’s encoder and decoder, resp.; and $K,~{}M,~{}T$ be the message-length, sub-message-length, and number of channel uses, resp., with $T_{M}=TM/K$. We report the Big-$\mathcal{O}$ order of FLOPS for each method in Tab. I, [^6] Full (handwritten) derivation located in our GitHub repository. For subsequent experiments, we set $h_{c}=32$, $h_{b}=32$, and $h_{r}=50$, leaving approximately $0.6M$, $2.5M$, and $2.6M$ FLOPS for TWLC, TWBAF, and TWRNN, respectively.
TABLE I: Big-$\mathcal{O}$ FLOPS comparison between models, where $K,~{}M,~{}T,~{}h_{\{c,b,r\}}$ are the message-length, sub-message-length, total number of channel uses, and hidden dimensions, resp. Equations are for a single user’s encoder and decoder, i.e., half of the total FLOPS.
@@
TWLC $\mathcal{O}(Th_{c}(M+T_{M})+h_{c}^{2})$ $\mathcal{O}(\nicefrac{{Kh_{c}}}{{M}}(M+T_{M})+h_{c}^{2}+h_{c}2^{M})$
TWBAF $\mathcal{O}(T_{M}(L_{B}^{E}\ell h_{B}^{2}+\ell h_{B}(T_{M}+M)))$ $\mathcal{O}(L_{B}^{D}\ell h_{B}^{2}+\ell h_{B}(M+T_{M})+2^{M+1}\ell h_{B})$
TWRNN $\mathcal{O}(Th_{r}^{2}+Mh_{r})$ $\mathcal{O}(\nicefrac{{K}}{{M}}(T_{M}(h_{r}^{2}+Mh_{r})+2^{M}h_{r}))$
@@ IV-B Comparison of One-way Versus Two-way Communications
In this section, we explore OW versus TW feedback communication. In Tab. II, ALC shows considerable BLER improvement over LC, making it more competitive for comparison. We consider two cases: TWLC versus ALC, and the TW linear scheme (TWLIN) from [16] against its predecessor, the inner code of Chance and Love (CL) [5]. [^7] We consider the CL scheme with symbols from a 4-PAM constellation. Users transmit $K$ bits apiece over a total of $T$ channel uses. For the OW case, each user is allotted $T/2$ channel uses, so $R_{ow}=2K/T$. For TW, the length-$K$ bitstreams are split into sub-blocks of length $M=K/2$; those $M$ bits are transmitted over $T/2$ channel uses, yielding $R_{tw}=K/T$. The decoders’ outputs are then concatenated for BLER calculation. We consider settings of $K=6,~{}T=18$ and $K=4,~{}T=16$.
Looking first to the ML-based models, we find in Fig. 3(a) distinct regions in which OW and TW communications are advantageous. For $K=6$, the lower coding rate enabled by TW coding admits uniformly lower BLER. However for $K=4$, the OW model generally performs better. Turning to the linear schemes, the $K=6$ data seems to reverse the BLER/SNR trend from TWLC/ALC. We may attribute these results to the optimization scheme of TWLIN, which we found not particularly robust. For $K=4$, TWLIN shows uniform improvement over CL, again showing benefit in the lower coding rate permitted by TW schemes. These results suggest that while there is great potential benefit in utilizing TW schemes, it is worthwhile to benchmark against a comparable OW scheme when deploying a model to ensure a desirable BLER/complexity tradeoff.
TABLE II: BLER performance of one-way Active LightCode vs standard LightCode for $SNR_{2}=20$ with $K=M=3$ and $T=9$.
@@
BLER (ALC) $69.8\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 6 $83.0\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 8 $94.0\mathrm{e}\hskip 1.13809pt\text{--}\hskip 1.13809pt$ – 9
@@ IV-C Two-Way BLER Comparison
Figure 3: Sum BLER results for all experiments. “TW” is truncated from the model names in legends in Figs. (b-f) for brevity. In (a), we compare OW and TW communications transmitting the same total number of bits over the same total channel uses (yielding rate $R$, as explained in Sec. IV-B), finding distinct regions over which each is more advantageous. (b) shows the sensitivity of each model to out-of-training-distribution noise, finding that matching noise between training and inference can play a substantial role in performance. (c,d) and (e,f) explore rate $1/3$ and $2/5$ codes resp., with $K$ denoting the total number of bits, $T$ the block-length, and $M$ the “sub-block” size into which a length-$K$ bitstream is split for encoding. “POL (OL)” denotes an open-loop TW channel with each user using a $(T,K)$ polar code.
In Fig. 3 (c-f), we explore rate $R$ codes on messages of length $K$. For length-$M$ sub-blocks, $TM/K$ channel uses apiece are utilized in transmitting $K/M$ sub-blocks whose results are concatenated for BLER calculation of the length-$K$ block. In Fig. 3(c), all models perform comparably in lower $SNR_{2}$ regimes, making TWLC, with its smaller architecture, a viable option. TWRNN and TWBAF perform noticeably better as $SNR_{2}$ increases, indicating their ability to capture more complex relationships between the users. For $SNR_{1}=-1$ dB in Fig. 3(d,f), we find that TWRNN clearly outperforms all other models. The performance trends seen in the rate $1/3$ setting continue into the rate $2/5$ regime of Fig. 3(e,f), with the TWLIN demonstrating mostly inferior BLER to the ML-based models. While we expect model performance would suffer at higher coding rates, this is not the job of these physical layer models; such higher coding rates are achieved via outer codes.
Lastly in Fig. 3(b), the sensitivity of each model is assessed against out-of-distribution (OOD) noise by first training a “fixed” model on $SNR_{1}=1~{}\text{dB},~{}SNR_{2}=15$ dB then testing the model as a function of $SNR_{2}$. While all models are reasonably robust to OOD noise, there is distinct benefit in tailoring the training to closely match the intended use-case, as evidenced by the superiority of the lines labeled “Orig”, run with matched train/test SNR. We also observe that TWLC trained under fixed $SNR_{2}$ has an error floor that is noticeably higher than the other models, reached around $SNR_{2}=15$ dB. This indicates that the more intricate RNN and attention mechanism structures employed by TWRNN and TWBAF may enable better generalizability to settings with favorable noise conditions during inference relative to training.
@@ IV-D Key Takeaways
TWLC, TWBAF, and TWRNN all demonstrate viability under different engineering goals and system settings. From Sec. IV-B, we find that one-way and two-way communications demonstrate distinct SNR regions and coding rates over which they respectively perform better. Examining the distinct models, we find that TWLC obtains decent performance with a significantly smaller network, requiring under $25\%$ of the FLOPS to run as the other two (Sec. IV-A). However, TWLC is relatively less capable of capturing the intricate patterns available in certain SNR regimes. TWRNN demonstrates superior BLER performance in low $SNR_{1}$ regimes but suffers from large size and computational cost due to its recurrent structure. Lastly, TWBAF performs comparably to TWRNN in higher $SNR_{1}$ settings, making it a viable alternative. The high performance of TWRNN and TWBAF appears in large part to be due to their utilization of attention mechanisms.
@@ V Conclusions
In this paper, we presented an algorithmic framework for ML-based two-way feedback coding. We then explored the performance tradeoffs of one-way and two-way communications. We find that in some noise regimes, two-way coding demonstrates superior performance to comparable one-way systems. Lastly, we explored the efficacy of three two-way models, finding that each is meritorious depending on the desired engineering outcomes. These efforts can guide future endeavors in the development of novel coding schemes, adaptive feedback mechanisms, and performance optimization.